GlueX TOF Meeting, January 27, 2021
GlueX Time-of-Flight Meeting
Wednesday, January 27, 2021
10:00 am EST
BlueJeans: 350 531 998
- Review of minutes from the last meeting
- PMT Repair
- TOF II Trigger
- TOF NIM paper
- Action Item Recap
Talks can be deposited in the directory
/group/halld/www/halldweb/html/talks/2021 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2021/ .
Present: Eugene Chudakov, Mark Ito (chair), Sasha Ostrovidov, Susan Schadmand, Beni Zihlmann
Review of minutes from the last meeting
We went over the minutes from January 13 without significant comment.
- Prompted by Sasha's question at the last meeting, Beni checked on the HV settings of the tubes that are candidates for replacement. He found that they are all on the low side, 1210 to 1290 V. Based on that, he thinks that we should replace only two of the PMTs (PMTs 69 and 113) rather than the five he identified last time.
- Sasha reminded us that all of the single-ended, new-in-TOF-II modules have the end surface near the beam hole blackened.
- Sasha told us that Paul Eugenio is willing to come up to the Lab to do the replacement, but the required two-week quarantine makes that impractical.
- The tubes will be replaced in situ (no need to remove the entire array from the forward carriage).
- Beni proposed practicing PMT replacement on one of the unused modules in the Hall.
- Mark suggested that as we learn how to do the replacement, we also train a member of the technical staff on the task.
TOF II Trigger
Beni ran down the activity in progress on standing up a trigger with TOF II.
- We want to present progress at the CPP/NPP readiness review.
- Hai Dong told Beni that all changes necessary to support the TOF II geometry can be done in the crate trigger processor (CTP). Nothing upstream of the CTP need be modified.
- Hai has written up a work plan. Elton presented it at the CPP meeting this morning.
- The work will take 4 weeks.
- 2 to 3 weeks will be needed for bench testing.
- Sascha Somov will have to test the system in the Hall after that, either with cosmics or (most likely) with beam.
- The plan puts us on track for having the system ready for beam tests before the August run.
- Beni and Mark will meet to confirm channel assignments as transmitted to Hai. Mark has the map that was used to create translation tables that have worked in recent data taking.
TOF NIM paper
No known progress.
Action Item Recap
- Decide on when to ship PMTs to JLab from FSU (Mark, Paul Eugenio)
- Prepare a complete draft of the TOF NIM paper. (all)
- Make a plan for TOF repair during the SAD, including Paul in the discussion. (all)
- Renew the equipment loan agreement between JLab and FSU (Mark, Paul)